OpenStudy (anonymous):

factor the trinomial 16+6x-x^2 Hint: (m-x)(n+x) (find m and n)

OpenStudy (anonymous):

=2*8 + (8-2)x-x^2 =(2+x)(8-x) m=8,n=2

OpenStudy (anonymous):

so how would I write this in simplest form

OpenStudy (anonymous):

(2+x)(8-x)